OpenStudy (rock_mit182):

(a) How much charge is on the surface of an isolated spherical conductor that has a radius 10cm and is charged to 2KV (b) What is the electrostatic potential energy of this conductor? (Assume the potential is zero far from the sphere.)

OpenStudy (rock_mit182):

@Michele_Laino

OpenStudy (rock_mit182):

I guess if the radius is infinite the potential is 0.. but im stuck from there

OpenStudy (michele_laino):

the relationship between the potential \(V\) on the surface of our sphere, and its charge is: \[V = \frac{1}{{4\pi {\varepsilon _0}}}\frac{Q}{a}\] where \(a\) is the radius of our sphere

OpenStudy (rock_mit182):

inside the sphere there will be constant potential, right?

OpenStudy (michele_laino):

yes! since inside the sphere there is no electric field

OpenStudy (michele_laino):

It is supposed that the electric charge is only on the surface of the sphere

OpenStudy (rock_mit182):

ok i got the first part 2* 10^3 V (4*pi*epsilon_o)(.1) m =Q

OpenStudy (rock_mit182):

what about b) ?

OpenStudy (michele_laino):

for part (b), we have to consider the energy trapped inside the electrostatic field outside the sphere

OpenStudy (rock_mit182):

that's why faraday's cage is a success ?

OpenStudy (michele_laino):

yes! inside a metallic box there is no electric field when, of course, outside that metallic box there is one or more than one electric field

OpenStudy (rock_mit182):

that's really awesome

OpenStudy (michele_laino):

yes! that behaviour it is due to the fact that the electrons of the metallic box will redistribute themselves such that they create an electric field which cancel the external electric field

OpenStudy (michele_laino):

for part (b), we have to compute this integral: \[\Large W = \int_a^{ + \infty } {\frac{1}{{8\pi }}} \frac{{{Q^2}}}{{{r^4}}}4\pi {r^2}dr\] where I have used the \(CGS\) system

OpenStudy (michele_laino):

since I assign a density energy of \[\frac{{{E^2}}}{{8\pi }}\] in the outer space

OpenStudy (rock_mit182):

ok now i get what is inside the integral

OpenStudy (rock_mit182):

However idk where did you get that density

OpenStudy (michele_laino):

of course we can show why we have that energy density( \(energy/Volume\) ), nevertheless, when I make computations similar to your exercise I simply remember the formula of that density

OpenStudy (michele_laino):

I remark that I have used \(CGS\) system, since when I was at university I have studied electromagnetism using that system

OpenStudy (rock_mit182):

mm I see, did you study physics i guess

OpenStudy (michele_laino):

yes! I have a degree in Physics the electric field, written in \(CGS\) system is: \[E = \frac{Q}{{{r^2}}}\]

OpenStudy (rock_mit182):

well I have another question, maybe is unrealated to this topic, Is there such thing as magnetic potential or electromagnetic potential ?

OpenStudy (michele_laino):

more precisely there is a vector potential \(A\) for magnetic field \(B\), such that the subsequent vector equation holds: \[{\mathbf{B}} = {\text{rot }}{\mathbf{{\rm A}}}\]

OpenStudy (michele_laino):

it is a partial differential equation, and \(rot\) stands for the \(rotor\) operator applied to the vector \(A\)

OpenStudy (rock_mit182):

i guess I have to learn multivariable calculus to understand such concepts...

OpenStudy (rock_mit182):

and of course even more differential equations ..

OpenStudy (michele_laino):

yes! correct! Since the vector \(A\) is a function of the spatial variable \(x,y,z\) and, for time-dependent fields, it is also a function of time \(t\)

OpenStudy (michele_laino):

variables*

OpenStudy (rock_mit182):

well dude you have been very helpful, now i have to study capacitors if a qeustion comes along I'll looking for you. thanks for helping me out

OpenStudy (michele_laino):

thanks! :) Finally, we can write this analogy for both electric and magnetic field: Electric field: \[{\mathbf{E}} = grad\varphi \] and, magnetic field: \[{\mathbf{B}} = {\text{rot }}{\mathbf{{\rm A}}}\] the functions \(\phi\) and \(A\) play analogue role

OpenStudy (michele_laino):

\(grad\) stands for \(gradient\) operator applied to function \(\phi\)

OpenStudy (rock_mit182):

gradient is a directional derivative, right ?

OpenStudy (michele_laino):

yes! In cartesian coordinates we have this formula: \[{\text{grad}}\varphi = \left( {\frac{{\partial \varphi }}{{\partial x}},\frac{{\partial \varphi }}{{\partial y}},\frac{{\partial \varphi }}{{\partial z}}} \right)\] it is a vector

OpenStudy (rock_mit182):

damn This is really nice sometimes math could be outstanding

OpenStudy (michele_laino):

of course! :)